OpenStudy (anonymous):

For which of the following function does e range contain both positive and negative values over the domain 3less then or equal to x less then or equal to 7? ( sorry i could do the sign so i just write it out hopefully u understand it ) I. f(x)= 3x-10 II. g(x)=2x-7/x-4 III. h(x)= 5x+9/x-6 A) I only B) I and II only C) I and III only D) II and III only E) I,II, and III

OpenStudy (anonymous):

domain is \[3\leq x \leq 7\] right?

OpenStudy (anonymous):

so for example the first one has range that contains positive and negative numbers because \[f(3)=-1\] and \[f(7)=11\]

OpenStudy (anonymous):

yup yup

OpenStudy (anonymous):

\[g(x)=\frac{2x-7}{x-4}\] certainly contains positive and negative values

OpenStudy (anonymous):

has a vertical asymptote at x = 4 and as you approach 4 from the left it goes to - infinity, from the right it goes to positive infinity, so it contains both positive and negative values

OpenStudy (anonymous):

exact same thing for the last one, so if i read the question correctly the answer is all three have positive and negative values. what kind of class is this for?

OpenStudy (anonymous):

SATclass

OpenStudy (anonymous):

yuck